40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8562     Entry time: Sat May 11 01:11:52 2013
Author: Koji 
Type: Update 
Category: ASC 
Subject: PRC mode stabilization with a shadow sensor at POP 

Ah, AWESOME. Indefinite PRMI lock was finally achieved.

POP setup

- Looked at the POP setup. Checked the spot on POP110 PD. Found some misalignment of the beam.
  The beam spot was aligned to the PD with PRMI locked. The value of POP110I almost doubled by the alignment
  and recovered previous value of 400. Therefore previous normalization values of MICH 0.01 / PRCL 100 were restored.

- Placed PDA36A (Si 3.6mmx3.6mm) on the POP path that Jenne prepared. The gain knob was set to 40dB.
  Since the original spot had been too small, a lens with f=50mm was inserted in order to expand the beam.
  Connected the PD output to the SMA feedthrough on the ITMX table enclosure.
  I found the BNC cable labeled "PO DC" hanging. Connected this cable to the enclosure SMA.

- Went to the LSC rack. Found the corresponding PO DC cable. Stole the POPDC channel from POP110I Bias T to this PO DC cable.

- Razor blade setup: Machined a junk Al bracket in order to fix a razor blade on it. Attached the Al bracket to a sliding stage.

Locking

- Locked the PRMI with REFL33I&AS55Q. Cut the beam into half by the razor blade.

- Made a temporary PRM_ASC_YAW filter.
  Zero: 0Hz Pole: 2kHz
  Resonant Gain 3.2Hz Q:2 Height 30dB
  Butterworth 2nd-order 60Hz

  => Expected UGF 0.1Hz&10Hz

- CDS: By the work described in this entry, the POPDC signal was connected to the "MC" bank of the LSC.
            BTW, the 11th row of the LSC output matrix is connected to the PRM_ASC_YAW.

- The "MC" servo input (i.e. the POPDC signal) was normalized by POP110I (without SQRTing).

- Engaged the PRM ASC path. Gradually increased the gain of PRM_ASC_YAW. G=+100 seemed to be the best so far.
  It was visible that the spot on the POP CCD was stablized in yaw.

- The lock lasted for ~40min. Took several measurements, alignment adjustment, etc.

- Tweaking the PRM ASC unlocked the PRMI.

- Locked again. Switched from REFL33I/AS55Q (x1/x1) combination to REFL55I/REFL55Q (x1/x0.3) combination.
  This also kept the lock more than 20min.

Attachment 1: Screenshot.png  48 kB  Uploaded Sat May 11 02:49:27 2013  | Hide | Hide all
Screenshot.png
Attachment 2: 130510_PRMI.pdf  81 kB  Uploaded Sat May 11 11:59:48 2013  | Hide | Hide all
130510_PRMI.pdf
ELOG V3.1.3-